# Dichography: Two-frame Ultrafast Imaging from a Single Diffraction Pattern

### v2

@@ Acknowledgments
Appendix A Algorithmic approach to Dichography
A.1 Iterative phase retrieval
@@
Figure 5: Scheme of the iterative phase retrieval algorithm for Dichography. The approach corresponds, in most steps, to the execution of two independent reconstruction procedures, AA and $B$, using conventional iterative phase retrieval algorithms. The two densities, $\rho^{A}$ and $\rho^{B}$, are randomly initialized. Two constraints – the support function in real space and the experimental intensities in Fourier space – are cyclically enforced. The application of the real-space constraint (shown on the right and left sides of the figure) is independent for AA and $B$, each with its own support function. This step is thus equivalent to conventional methods for two independent reconstructions. The Fourier intensity constraint, depicted in the center of the figure, is shared between AA and $B$. Here, the sum of the two Fourier amplitudes is replaced by the experimentally measured values. This information is then back-propagated to the two separate densities, $\rho^{A}$ and $\rho^{B}$. The two reconstruction outcomes are returned once the total number of iterations is reached. For sake of clarity, the knowledge of the correct support functions is assumed a-priori. However, in all reconstructions presented in this work, the supports are unknown and recovered along the reconstruction process, as discussed in Appendix A.3.
@@
Coherent Diffraction Imaging (CDI) is an indirect imaging method that requires extensive numerical analysis to obtain the sample image from the recorded data. The scattered field $\psi$ encodes the complete information about the spatial density of the sample, $\rho$, as they are mathematically linked by a Fourier Transform (FT) operation (under certain approximations [7]). It is therefore possible to reconstruct $\rho$ from the full knowledge of the field, i.e., $\rho\propto\mathcal{F}^{-1}\left[\psi\right]$ where $\mathcal{F}^{-1}$ indicates the inverse FT operation. However, the experimentally recorded image, $I$, provides information only on the field’s amplitude, i.e., $I\propto\left|\psi\right|^{2}$, while the field’s phase is lost. Recovering this lost phase corresponds to reconstructing the sample density $\rho$. This so-called phase retrieval problem lies at the core of CDI analysis and must be solved by suitable numerical approaches.
@@
Conventional algorithms for phase retrieval, such as the Hybrid Input-Output (HIO) and Error Reduction (ER) methods [59], are characterized by an iterative scheme, as they cyclically impose two constraints [32]. On one hand, the oversampling condition is enforced by setting to zero the density values outside the spatial extension of the sample. This is achieved by defining a support function, a binary-valued matrix whose entries are set to $1$ where non-zero density of the sample is expected, and to 0 where zero scattering is enforced. On the other hand, the intensity constraint is applied in reciprocal space, where the experimentally measured intensities replace the Fourier amplitudes of the current density estimate. This iterative approach is retained for Dichography and lies at the core of the numerical retrieval of the two frames, $\rho^{A}$ and $\rho^{B}$.
@@
Fig. 5 shows a flowchart of the iterative phase retrieval algorithm adapted to Dichography. It comprises two reconstruction workflows, one for each frame, visually separated on the left and right sides of Fig. 5. Each workflow follows the typical scheme of conventional iterative algorithms [59, 32]. Both densities, $\rho^{A}$ and $\rho^{B}$, are independently initialized, typically with random values, and each is constrained to its own support function in real space, $\textsc{s}^{A}$ and $\textsc{s}^{B}$, respectively. However, unlike conventional CDI, the two procedures are entangled when constraining the experimental intensities in Fourier space. This step is shown at the center of Fig. 5.
@@
A.2 Constraining experimental data
In Dichography, the experimentally acquired information on the sum of the two scattering signals has to be constrained, i.e., propagated to the two frames of the reconstruction, to make the two estimates of the field, $\psi^{A}$ and $\psi^{B}$, satisfy the condition in Eq. (2).
@@ A.2 Constraining experimental data
Unlike conventional CDI – in which only a single phase needs to be recovered for each field amplitude – the phase retrieval problem in Dichography is three-fold. Indeed, it requires the recovery of three phase values: $\phi^{A}$ and $\phi^{B}$ for the two scattered fields, as well as $\phi^{M}$, which encodes the relative fields amplitude.
Differently from conventional CDI – where for each field amplitude only as single phase has to be recovered –, the phase retrieval problem in Dichography is three-fold, requiring the retrieval of three phase values, $\phi^{A}$ and $\phi^{B}$ for the two scattered fields and $\phi^{M}$ that encodes their relative amplitude.
A detailed and formal presentation of the mathematical operations involved is provided in Sec. S3 of the Supplemental Material [34].
The most common iterative phase retrieval algorithms share the same way of constraining Fourier intensities, while they differ in the application of the support constraint. The modified intensity constraint described in Fig. 6 can thus be directly implemented in all conventional iterative phase retrieval algorithms [32, 59]. The behavior of the algorithms, however, differs from their conventional CDI counterparts due to the different mathematical nature of the underlying optimization problem (as discussed in Sec. S4 of the Supplemental Material [34]).
Conceptually similar implementations of this intensity constraint have been developed and employed to deal with broadband diffraction data, in cases where the spectral information is known [60, 61], or even where the spectrum of the scattered light is retrieved along the reconstruction process [62]. These approaches have opened new possibilities for CDI with broadband or polychromatic pulses, like those delivered by High-Harmonic Generation sources [63]. Still, differently from those methods, Dichography doesn’t enforce any structural connection between the contributions to the recorded scattering signals. This allows capturing structural changes in the samples, as discussed in the following section.
@@
A.3 Real-space constraint
@@
The application of the real-space constraint, i.e., the support function, is performed independently on the two frames, as if they were two separate reconstruction procedures using iterative algorithms for conventional single-particle CDI. This means that each of the two densities, $\rho^{A}$ and $\rho^{B}$, is constrained to its own independent support function, $\textsc{s}^{A}$ and $\textsc{s}^{B}$, respectively, as shown in Fig. 5. The way the support constraint is enforced depends on the specific iterative algorithm employed [32] and retains its form in the adaptation for Dichography.

### v3

@@ Acknowledgments
At the time of the original submission, the authors were not aware of the theoretical and algorithmic work of Millane and Chen [33]. We thank Prof. Richard Kirian (Arizona State University, USA) for bringing this work to our attention. We are especially grateful to Prof. Rick Millane (University of Canterbury, New Zealand) for the feedback and suggestions on the original version of the manuscript, which significantly contributed to the improved quality of the theoretical arguments presented in this revised version.
@@
V Methods
V.1 Algorithmic approach to Dichography
V.1.1 Iterative phase retrieval
@@
Figure 5: Scheme of the iterative phase retrieval algorithm for Dichography. The approach corresponds, in most steps, to the execution of two independent reconstruction procedures, AA and $B$, using conventional iterative phase retrieval algorithms. The two densities, $\rho^{A}$ and $\rho^{B}$, are randomly initialized. Two constraints – the support function in real space and the experimental intensities in Fourier space – are cyclically enforced. The application of the real-space constraint (shown on the right and left sides of the figure) is independent for AA and $B$, each with its own support function. This step is thus equivalent to conventional methods for two independent reconstructions. The Fourier intensity constraint, depicted in the center of the figure, is shared between AA and $B$. Here, the sum of the two Fourier amplitudes is replaced by the experimentally measured values. This information is then back-propagated to the two separate densities, $\rho^{A}$ and $\rho^{B}$. The two reconstruction outcomes are returned once the total number of iterations is reached. The support function is periodically updated and refined, based on the current density estimate, completely independently between the two frames, as discussed in Sec. V.1.3
@@
Coherent Diffraction Imaging (CDI) is an indirect imaging method that requires extensive numerical analysis to obtain the sample image from the recorded data. The scattered field $\psi$ encodes the complete information about the spatial density of the sample $\rho$ as they are mathematically linked by a Fourier Transform (FT) operation (under certain approximations [7]). It is therefore possible to reconstruct $\rho$ from the full knowledge of the field, i.e., $\rho\propto\mathcal{F}^{-1}\left[\psi\right]$ where $\mathcal{F}^{-1}$ indicates the inverse FT operation. However, the experimentally recorded image, $I$, provides information only on the field’s amplitude, i.e., $I\propto\left|\psi\right|^{2}$, while the field’s phase is lost. Recovering this lost phase corresponds to reconstructing the sample density $\rho$. This so-called phase retrieval problem [64] lies at the core of CDI analysis and must be solved by suitable numerical approaches.
@@
Conventional algorithms for phase retrieval, such as the Hybrid Input-Output (HIO) and Error Reduction (ER) methods [36], are characterized by an iterative scheme, as they cyclically impose two constraints [34, 37, 65]. On the one hand, the oversampling condition is enforced by setting to zero the density values outside the spatial extension of the sample. This is achieved by defining a support function, a binary-valued matrix whose entries are set to $1$ where non-zero density of the sample is expected, and to 0 where zero scattering is enforced. On the other hand, the intensity constraint is applied in reciprocal space, where the experimentally measured intensities replace the Fourier amplitudes of the current density estimate. This iterative approach is retained for Dichography and lies at the core of the numerical retrieval of the two frames, $\rho^{A}$ and $\rho^{B}$.
@@
Fig. 5 shows a flowchart of the iterative phase retrieval algorithm adapted to Dichography. It comprises two reconstruction workflows, one for each frame, visually separated on the left and right sides of Fig. 5. Each workflow follows the typical scheme of iterative projections algorithms [36, 34, 37]. Both densities, $\rho^{A}$ and $\rho^{B}$, are independently initialized, typically with random values, and each is constrained to its own support function in real space, $\textsc{s}^{A}$ and $\textsc{s}^{B}$, respectively. However, unlike conventional CDI, the two procedures are entangled when constraining the experimental intensities in Fourier space. This step is shown at the center of Fig. 5.
@@
V.1.2 Constraining experimental data
In Dichography, the experimentally acquired information on the sum of the two scattering signals has to be constrained, i.e., propagated to the two frames of the reconstruction, to make the two estimates of the field, $\psi^{A}$ and $\psi^{B}$, satisfy the condition in Eq. (2). Its operation is equivalent to the one proposed in Ref. [33] for the specific case of two incoherent contributions to the scattering.
@@ V.1.2 Constraining experimental data
Unlike conventional CDI – in which only a single phase needs to be recovered for each field amplitude – the phase retrieval problem in Dichography is three-fold. Indeed, it requires the recovery of three phase values: $\phi^{A}$ and $\phi^{B}$ for the two scattered fields, as well as the relative fields amplitude encoded in the “fictitious” phase $\phi^{M}$. However, this observation doesn’t imply that the dichographic problem is three times harder, thanks to the action of the support constraints [33] as further discussed in Sec. V.2.1. A more detailed presentation of the mathematical operations involved is provided in Sec. S3 of the Supplemental Material [39], and thoroughly discussed and extended to an arbitrary number of incoherent scattering contributions in Ref. [33].
The most common iterative phase retrieval algorithms share the same way of constraining Fourier intensities, while they differ in the application of the support constraint. The modified intensity constraint described in Fig. 6 can thus be directly implemented in all conventional iterative phase retrieval algorithms [34, 36, 37], as shown in the case of the Difference Map [33]. The behavior of the algorithms, however, differs from their conventional CDI counterparts due to the different mathematical nature of the underlying optimization problem (as discussed in Sec. S4 of the Supplemental Material [39]).
Conceptually similar implementations of this intensity constraint have been developed and employed to deal with broadband diffraction data, in cases where the spectral information is known [66, 67], or even where the spectrum of the scattered light is retrieved along the reconstruction process [68]. These approaches have opened new possibilities for CDI with broadband or polychromatic pulses, like those delivered by High-Harmonic Generation sources [69]. Crucially, unlike those methods, Dichography doesn’t enforce any structural connection between the contributions to the recorded scattering signals and doesn’t involve any a priori knowledge on the sample’s shape. This allows capturing structural changes in the samples, as discussed in the following section.
@@
V.1.3 Real-space constraint
@@
The application of the real-space constraint, i.e., the support function, is performed independently on the two frames, as if they were two separate reconstruction procedures using iterative algorithms for conventional single-particle CDI. This means that each of the two densities, $\rho^{A}$ and $\rho^{B}$, is constrained to its own independent support function, $\textsc{s}^{A}$ and $\textsc{s}^{B}$, respectively, as shown in Fig. 5. The way the support constraint is enforced depends on the specific iterative algorithm employed [34, 37] and retains its form in the adaptation for Dichography.